OpenStudy (anonymous):

identify the oblique asymtote of f(x)=x^2+6x-9/x-2

OpenStudy (anonymous):

no oblique asymptote can be found but found a vertical one :D

OpenStudy (amistre64):

P(x) = Q(x) + R(x), such that asymptote = Q(x)

OpenStudy (amistre64):

divide it thru and toss your remainder

OpenStudy (amistre64):

x+8 <-- asymptote ------------ x-2 | x^2+6x-9 -x^2+2x ------- 8x -9 -8x+16 --------- 7 , toss the 7

OpenStudy (anonymous):

Thank you!!! :)

OpenStudy (amistre64):

youre welcome

myininaya (myininaya):

toss the 7 where?

myininaya (myininaya):

i would put it in the garbage can i wouldn't litter like amistre would

OpenStudy (amistre64):

7 up?

OpenStudy (anonymous):

I was a bit confused about the toss the 7 part.

OpenStudy (amistre64):

the 7 part is a remainder that can be placed in at the end like this: \[\frac{7}{x-2}\] now, when x gets very large ... we are going towards infinity here .... this part goes eventually gets so small that it equals 0

OpenStudy (amistre64):

so we can ignore it as part of the asymptote

OpenStudy (anonymous):

so would the answer be -8x+16/7(x-2)?

OpenStudy (amistre64):

no, the asymptote line that it approaches is: x+8 when x = infinity, then we can talk about it touching the asymptote

OpenStudy (amistre64):

think of an asymptote as a line that the graph snuggles up to, but never touches

OpenStudy (amistre64):

it gets closer and closer, but never hits it unless we can find an end to infinity

OpenStudy (amistre64):

in other words, out there in the far off distance, that remainder part is useless and disappears into eternity